OpenStudy (anonymous):

Find the center and radius of the circle whose equation is x^2+2x+y^2+4y-11=0

OpenStudy (dumbcow):

you want to end up in this form: (x-h)^2 + (y-k)^2 = r^2 have you used completing the square? that is how you do this problem

OpenStudy (anonymous):

I got the center as (2,1) and the radius as (-6,2)

OpenStudy (dumbcow):

hmm well radius is a length not a point and the center is incorrect

OpenStudy (anonymous):

I got this vertex now 2, -6

OpenStudy (dumbcow):

how are you getting this? if you dont show steps i cant know what you are doing wrong also vertex if for parabolas ..... circles dont have vertex, only center and radius

OpenStudy (dumbcow):

completing the square (x^2 +2x + ?) = (x+a)^2 --> (x+a)^2 = x^2 + 2ax + a^2 --> x^2 +2x +? = x^2 +2ax +a^2 therefore a = 1 and a^2 = 1 (x+1)^2 - 1= (x^2 +2x)

OpenStudy (dumbcow):

do same thing for y (y^2 +4y) = (y+2)^2 -4 now you have (x+1)^2 -1 +(y+2)^2 -4 = 11 (x+1)^2 + (y+2)^2 = 16 center: (-1,-2) radius = sqrt16 = 4